A lab tech has one solution that is 45% acid and another that is 70 % acid. How much of each is needed to make 100 mL of a 65% acid solution?
Answer by stanbon(75887)

A lab tech has one solution that is 45% acid and another that is 70 % acid.
How much of each is needed to make 100 mL of a 65% acid solution?
----
Equation:
acid + acid = acid
0.45x + 0.70(100-x) = 0.65*100
-----
Multiply thru by 100 to get:
45x + 70*100 - 70x = 65*100
5*100 = 25x
x = 20 mL (amt of 45% solution needed)
100-x = 80 mL (amt of 70% solution needed)
============================================
cheers,
Stan H.
=================